# TAVP: Task-Adaptive Visual Prompt for Cross-domain Few-shot Segmentation

### v1

@@ Abstract
Under the backdrop of large-scale pre-training, large visual models (LVM) have demonstrated significant potential in image understanding. The recent emergence of the Segment Anything Model (SAM) has brought a qualitative shift in the field of image segmentation, supporting flexible interactive cues and strong learning capabilities. However, its performance often falls short in cross-domain and few-shot applications. Transferring prior knowledge from foundation models to new applications while preserving learning capabilities is worth exploring. This work proposes a task-adaptive prompt framework based on SAM, a new paradigm for Cross-dominan few-shot segmentation (CD-FSS). First, a Multi-level Feature Fusion (MFF) was used for integrated feature extraction. Besides, an additional Class Domain Task-Adaptive Auto-Prompt (CDTAP) module was combined with the segmentation branch for class-domain agnostic feature extraction and high-quality learnable prompt production. This significant advancement uses a unique generative approach to prompts alongside a comprehensive model structure and specialized prototype computation. While ensuring that the prior knowledge of SAM is not discarded, the new branch disentangles category and domain information through prototypes, guiding it in adapting the CD-FSS. We have achieved the best results on three benchmarks compared to the recent state-of-the-art (SOTA) methods. Comprehensive experiments showed that after task-specific and weighted guidance, the abundant feature information of SAM can be better learned for CD-FSS.
@@
Cross-domain, Few-shot, Semantic Segmentation, Segment Anything
@@ I Introduction
Traditional deep networks have relied heavily on extensive annotated data to achieve high precision performance [1]. However, data annotation is a time-consuming task that requires significant human resources, especially for the segmentation of intensive pixel-level annotation tasks like medical images and remote sensing images. Therefore, Few-shot semantic segmentation (FSS) was introduced to narrow this gap [2], aiming to reduce the need for labeling. Besides, most of the few-shot learning (FL) methods mainly focus on learning the relationship between support and query sets within the same domain, which always requires fine-tuning on the target domain. The features extracted on the high level are class-agnostic but lack domain generalization, meaning that FL methods have the ability for new category generation but perform worse on new domain generation. The previous methods have limitations in that input-space-based enhancement requires expert knowledge to design enhancement functions, while feature-based enhancement usually relies on complex adversarial training [3]. Thus, Cross-domain Few-shot Segmentation (CD-FSS) [4] came up for solving segmentation tasks on medical, remote sensing, and other images, and four typical cross-domain data sets are illustrated in Figure 1.
Figure 1: The existing pipeline of cross-domain few-shot segmentation task.
@@
However, previous deep models may lead to poor generalization on unseen out-of-domain data, which limits their use in Cross-domain Few-shot scenarios. Recently, Large fundamental visual models (LVM) have made significant progress in natural image segmentation [5, 6, 7], including medical [8] and remote sensing [9] image segmentation. The Segment Anything Model (SAM) [10] was trained with over one billion masks and achieved unprecedented generalization capabilities on natural images. Additionally, some research has shown that proper adjustments to SAM can be applied in medical image segmentation [8] and zero-shot tasks. These advances suggest that powerful segmentation models with generalization capabilities can be used without designing complex networks as fine-tuning models due to the time-consuming retraining. Some early works have used pre-trained models on natural or medical images and achieved good performance [10, 11]. However, due to the inflexible capacity of pre-trained models and extensive few-shot methods such as disentanglement domain classifier [12], the cross-domain generalization ability of deep models has not been effectively improved.
@@
Although these LVM-based methods have improved the model’s performance in some professional domains, SAM does not have the ability to generalize in other vertical domains and requires a large number of vertical domain samples, which brings high expenses in data collection, sample labeling, and model training. At the same time, the vertical domain itself is difficult to exhaustively enumerate. Many works have integrated the SAM framework with meta-learning for transfer learning tasks [13, 14]. However, previous methods mainly focus on fine-tuning the SAM encoder, designing teacher-student by knowledge distillation [15], and feature matching by computing the distance between pairs. To increase computing efficiency and to better disentangle features of class and domains for model robust learning ability, we propose a Task-adaptive Visual Prompt (TAVP) algorithm achieving both inter and intra-domain information disentanglement combined with SAM. In contrast, our method can be effectively generalized to different vertical domains and can achieve results comparable to vertical domains. The pipeline comparison is shown in Figure 2.
Figure 2: (a)The existing cross-domain segmentation method based on SAM. (b) Our CDTAP module. (c)Produce class-specific and domain-specific feature. (d)Produce class-agnostic feature. (e)Produce domain-agnostic feature. (f)Original feature distribution.
@@
Upon further analysis of SAM, we identify its poor performance in CD-FSS, which can be attributed to a few key issues. The encoder’s image features, though containing basic class data, are mismatched with the target domain’s categories, with their inherent distributions potentially causing noise and performance drops. Effective learning requires feature information aligned with the target domain. Moreover, the decoder’s reliance on prompt-based cross-attention mechanisms also hinders its segmentation effectiveness.
@@
Based on the above analysis, the CDTAP module is proposed to better extract class and domain-specific features through contrastive learning from foreground and background for improving the robustness of CD-FSS, as shown in Figure 3. The experimental results show that our work can compute more accurate and robust pairing relationships between samples. Moreover, we propose a fully automatic segmentation framework based on SAM for CD-FSS. This new framework aims to enhance the model’s adaptability and accuracy for CD-FSS. Our contributions can be summarized as follows.
@@
  • Compared with SAM, which extracts features in multi-level context, we remain low-level feature representation and capture both global and local information by fusion.
@@
  • To realize the disentanglement of class and domain information, we integrate a unified and comprehensive feature transform method. Specifically, an additional Class-Domain Task-adaptive Auto-prompt (CDTAP) module is proposed for domain and class-agnostic feature extraction. Simultaneously, We use contrastive learning to achieve deeper and closer matching of samples among different domains.
  • To overcome the shortage of SAM that highly relies on human interaction, we propose an automatic and learnable prompt branch for segmentation tasks as weighted guiding and robust prompt generation.
  • We fine-tune SAM with less time and GPU consumption, and a heterogenization sampling strategy is adopted for Task-adaptive guiding segmentation.
  • Our method achieves competitive and best results on three benchmarks compared to the state-of-the-art methods on four publicly available CD-FSS data sets.
@@
The overall structure of the paper includes five sections. We briefly list some related technologies in cross-domain and few-shot learning in Section I. The review of related work about cross-domain and few-shot learning is described in Section II. We describe our method in more detail and show the model performance in Section III. Then, we conduct comprehensive experiments comparing previous methods and the SAM baseline in Section IV and  V. Finally, we conclude this paper and discuss the future application prospects in large model-based fine-tune methods in Section VI.
@@ II Related Works
We start this section by introducing a cross-domain segmentation task with its relative technology, and the few-shot segmentation task is described for related background. Then, we develop the CDFSS task and the related research within this field.
@@ II-A Domain Adaptation in Segmentation
In recent years, domain adaptation semantic segmentation has made notable progress. Domain adversarial training is utilized to learn domain-invariant representations in features [16]. Hoffman et al. [17] integrated global and local alignment methods with adversarial training. There were also other methods such as distillation loss [18], output space alignment [19], class-balanced self-training [20], and conservative loss [21], proposed according to a predefined curriculum learning strategy [22]. These methods collectively contributed to advancing adaptive semantic segmentation by leveraging information from various domains, ensuring the model’s robust performance across diverse and less annotated environments. Suppose the training data originate solely from a single domain and adaptation occurs to an unseen domain. Then, in this case, single-source domain adaptation becomes more challenging due to the limited diversity within the training domain. Consequently, a prevalent approach to address this issue is using data augmentation techniques to generate new domains, thereby enhancing the diversity and information content of the training data. Several methods were designed with varying generation strategies to address the single-source domain adaptation problem in computer vision tasks. For example, RandConv [23] employed random convolutions for data augmentation. MixStyle integrated style information from instances of randomly selected different domains.
In contrast with the above data augmentation methods, we use the foundation model to ensure rich prior knowledge instead of generating many images in the source domain, saving computing resources and increasing computation efficiency.
@@ II-B Few Shot Segmentation
The objective of Few-Shot Segmentation (FSS) tasks is to segment new semantic objects through a limited number of available labeled images or unlabeled images that are semantically distinct. Current methods primarily focused on improvements during the meta-learning phase. Prototype-based methods [24, 25, 26] utilized a technique where representative foreground or background prototypes were extracted from the support data, and various strategies were employed to interact between different prototypes or between prototypes and query features. Relation-based methods [27, 28, 29] also succeeded in few-shot segmentation. HSNet [30] built a high correlation using multi-scale dense matching and captures contextual information using 4D convolution. RePRI [31] introduced transductive inference of base class feature extraction that did not require meta-learning. However, these methods primarily focused on segmenting new categories from the same domain. Due to the significant differences in cross-domain distributions, they failed to be extended to unseen domains.
@@ II-C Cross-domain Few-shot Segmentation
There are four benchmarks [4] available for CDFSS standard evaluation. Figure  1 shows the difference between the original domain and CDFSS data sets. It is obvious that for the ChestX data set, the image format has been changed from RGB to gray, with a large gap from the original domain. The other two data sets have more edge information requiring high-quality semantic segmentation. RD [32] introduced a novel domain enhancement strategy leveraging a memory mechanism. This approach involved continuously storing domain-style information from the source domain during the training phase. Subsequently, during testing, this stored source information was utilized to enhance the segmentation performance. During testing, source domain information stored in memory was loaded for the target domain feature enhancement. RD [32] offered a direct approach to reduce domain differences and was validated on typical partitioned datasets. For semantic segmentation tasks in autonomous driving applications, PixDA [33] introduced an innovative pixel-by-pixel domain adversarial loss based on three key criteria: (i) aligning the source and target domains for each pixel, (ii) preventing negative transfer on correctly represented pixels, and (iii) regularizing the training of infrequent classes to mitigate overfitting. CDTF [34] achieved cross-domain few-shot segmentation by aligning support and query prototypes. This alignment was realized using an uncertainty-aware contrastive loss and supplemented with a supervised cross-entropy loss and an unsupervised boundary loss as regularization terms. CDTF [34] enabled the model to generalize from the base model to the target domain without requiring additional labels. CDFSS [35] presented a cross-domain few-shot segmentation framework that leveraged learning from natural domains to assist in rare-disease skin lesion segmentation. This approach was particularly valuable when dealing with limited data for common diseases in the target domain.
@@
Putting aside the previous simple computing prototype methods, we combine high-level prototype representation with a foundation model, SAM. Besides, we propose a dual prototype matching method for the foreground and background, ensuring fine-grained feature representation.
III Proposed Method
While SAM can be generalized to more scenarios, even zero-shot situations, it still has some limitations. Firstly, the original SAM relies on interactive prompts for accurate segmentation in different situations, which can be time-consuming. The second problem is how to transfer more abundant knowledge and key information from LVM methods while maintaining good generalization ability. To address the above two problems, we propose an automatic framework for segmentation with automatic prompts instead of interactive prompts by users. Meanwhile, an additional branch is designed for class and domain-agnostic feature extraction and Task-adaptive prompts generation.
The overall framework for TAVP is shown in Figure 3. The inputs of support and query from the source domain are fed into the SAM encoder for basic feature extraction. Note that we propose a multi-level feature fusion for extensive representation. Meanwhile, one target domain data is fed into the CDTAP module for class and domain specific and agnostic feature extraction. At the same time, this module generates learnable prompts as dense embedding input to the decoder. Then, the combined multi-level and dense prompts are fed into the SAM decoder for prediction.

### v2

@@ Abstract
While large visual models (LVM) demonstrated significant potential in image understanding, due to the application of large-scale pre-training, the Segment Anything Model (SAM) has also achieved great success in the field of image segmentation, supporting flexible interactive cues and strong learning capabilities. However, SAM’s performance often falls short in cross-domain and few-shot applications. Previous work has performed poorly in transferring prior knowledge from base models to new applications. To tackle this issue, we propose a task-adaptive auto-visual prompt framework, a new paradigm for Cross-dominan Few-shot segmentation (CD-FSS). First, a Multi-level Feature Fusion (MFF) was used for integrated feature extraction as prior knowledge. Besides, we incorporate a Class Domain Task-Adaptive Auto-Prompt (CDTAP) module to enable class-domain agnostic feature extraction and generate high-quality, learnable visual prompts. This significant advancement uses a unique generative approach to prompts alongside a comprehensive model structure and specialized prototype computation. While ensuring that the prior knowledge of SAM is not discarded, the new branch disentangles category and domain information through prototypes, guiding it in adapting the CD-FSS. Comprehensive experiments across four cross-domain datasets demonstrate that our model outperforms the state-of-the-art CD-FSS approach, achieving an average accuracy improvement of 1.3% in the 1-shot setting and 11.76% in the 5-shot setting.
@@
Cross-domain, Few-shot, Semantic Segmentation, Visual prompt
@@ I Introduction
Traditional deep networks relied heavily on extensive annotated data to achieve high precision performance [1]. However, data annotation is a time-consuming task that requires substantial human resources, particularly for intensive pixel-level annotation tasks such as medical image and remote sensing image segmentation. Therefore, Few-shot semantic segmentation (FSS) was introduced to narrow this gap [2], aiming to reduce the need for labelling. Besides, most of the few-shot learning (FL) methods mainly focus on learning the relationship between support and query within the same domain, which always requires fine-tuning on the target domain [3, 4]. The high-level features extracted are class-agnostic but lack domain generalization, meaning that while FL methods are capable of generating new categories, they perform poorly in generating new domains. The previous methods have limitations in that input-space-based enhancement requires expert knowledge to design enhancement functions, while feature-based enhancement usually relies on complex adversarial training [5]. Thus, Cross-domain Few-shot Segmentation (CD-FSS) [6] came up for solving segmentation tasks on medical, remote sensing, and other images.
@@
However, previous deep models may lead to poor generalization on unseen out-of-domain data, which limits their use in Cross-domain Few-shot scenarios. Recently, Large fundamental visual models (LVM) have made significant progress in natural image segmentation [7, 8, 9], including medical [10] and remote sensing [11] image segmentation. The Segment Anything Model (SAM) [12] was trained with over one billion masks and achieved unprecedented generalization capabilities on natural images. Additionally, some research has shown that proper adjustments to SAM can be applied in medical image segmentation [10] and zero-shot tasks. These advances suggest that powerful segmentation models with generalization capabilities can be used without designing complex networks due to time-consuming retraining. Some early works have used pre-trained models on natural or medical images and achieved good performance [12, 13]. However, due to the inflexible capacity of pre-trained models and extensive few-shot methods such as disentanglement domain classifier [14], the cross-domain generalization ability of deep models has not been effectively improved.
@@
While these LVM-based methods have enhanced model performance in certain professional domains, several limitations remain, which can be summarized as follows: (1) Poor Generalization to Specific Domains: While LVM-based methods enhance performance in certain professional fields, SAM struggles to generalize effectively to domain-specific tasks. (2) High Cost of Domain-Specific Adaptation: Adapting SAM to specific domains involves significant costs, including data collection, sample labelling, and model training. This dependency on large-scale domain-specific datasets makes the approach resource-intensive and impractical for domains with limited annotated data. (3) Difficulty in Covering All Domains: It is infeasible to exhaustively enumerate and address all possible specific domains. This inherent limitation restricts the scalability of SAM-based approaches in scenarios where diverse and evolving domain-specific needs are present. (4) Suboptimal Strategies for Transfer Learning: Although recent works have combined SAM with meta-learning for transfer learning, these methods [15, 16] primarily focus on fine-tuning the SAM encoder, implementing teacher-student frameworks through knowledge distillation, or using feature matching by pairwise distance computation [17]. These approaches often fail to provide comprehensive solutions for efficiently adapting SAM across diverse domains. To increase computing efficiency and to better disentangle features of class and domains for model robust learning ability, We propose a task-adaptive visual prompt (TAVP) algorithm that achieves both inter- and intra-domain information disentanglement. In contrast, our method can be effectively generalized to different vertical domains and achieves results comparable to state-of-the-art performance in those domains. The pipeline comparison is shown in Figure 1.
Figure 1: (a) The existing cross-domain segmentation method based on SAM. (b) Our Task-adaptive Visual Prompt. (i) There is no interaction between features from different categories in the original feature distribution. (ii) In a specific domain, prototypes are used to make semantic distinctions between categories to achieve clustering. (iii) Finally, inter-category distinction and intra-category strong constraints are achieved in a unified space.
@@
Upon further analysis of SAM, we observe its poor performance in CD-FSS, which can be attributed to a few key issues. The encoder’s image features, though containing basic class data, are mismatched with the target domain’s categories, with their inherent distributions potentially causing noise and performance drops. Effective learning requires alignment of feature information with the target domain. Moreover, the decoder’s reliance on prompt-based cross-attention mechanisms also hinders its segmentation effectiveness.
@@
Based on the above analysis, we proposed a CDTAP module to better extract class and domain-specific features through contrastive learning from foreground and background, to improve the robustness of CD-FSS, as shown in Figure 2. The experimental results show that our work can compute more accurate and robust pairing relationships between samples. Moreover, we propose a fully automatic segmentation framework based on SAM for CD-FSS. This new framework aims to enhance the model’s adaptability and accuracy for CD-FSS. Our contributions can be summarized as follows.
  • We propose a novel model that including a Multi-level Feature Fusion (MFF) and a Class Domain Task-Adaptive Auto-Prompt (CDTAP) module for efficiently combing SAM with CD-FSS task.
@@
  • Compared with SAM, which extracts features in a high-level context, the MFF is proposed to retain low-level feature representations and fuse global and local information to produce class-agnostic features.
@@
  • To realize the disentanglement of class and domain information, we integrate a unified and comprehensive feature transform method. Specifically, an additional Class-Domain Task-adaptive Auto-prompt (CDTAP) module is proposed for domain and class-specific feature extraction. Simultaneously, We use contrastive learning to achieve deeper and closer matching of samples among different domains.
  • To overcome the shortage of SAM that highly relies on human interaction, we propose an automatic, learnable prompt branch for segmentation, fine-tuning it efficiently with less time and GPU usage, and achieving competitive and best results compared to the state-of-the-art methods on four CD-FSS benchmarks.
@@
The overall structure of the paper includes five sections. We briefly list some related technologies in cross-domain and few-shot learning in Section I. The review of related work about cross-domain and few-shot learning is described in Section II. We describe our method in more detail and show the model performance in Section III. Then, we conduct comprehensive experiments comparing previous methods and the SAM baseline in Section IV. Finally, we conclude this paper and discuss the future application prospects in large model-based fine-tune methods in Section V.
@@ II Related Works
We start this section by introducing a cross-domain segmentation task with its relative technology, and the few-shot segmentation task is described for related background. Then, we develop the CD-FSS task and the related research within this field.
@@ II-A Domain Adaptation in Segmentation
In recent years, domain adaptation semantic segmentation has made notable progress. To enhance the domain adaptation method, CRTL was proposed by Wang et al. [18] by leveraging class priors and a projected Hilbert-Schmidt Independence Criterion (pHSIC) through transfer learning. Domain adversarial training is utilized to learn domain-invariant representations in features [19]. Hoffman et al. [20] integrated global and local alignment methods with adversarial training. Other approaches in domain adaptation have also been proposed, such as distillation loss [21], output space alignment [22], class-balanced self-training [23], and conservative loss [24], based on a predefined curriculum learning strategy [25]. These methods collectively contributed to advancing adaptive semantic segmentation by leveraging information from various domains, ensuring the model’s robust performance across diverse and less annotated environments. Suppose that the training data originate solely from a single domain and adaptation occurs to an unseen domain. Then, in this case, single-source domain adaptation becomes more challenging due to the limited diversity within the training domain. Consequently, a prevalent approach to address this issue is using data augmentation techniques to generate new domains, thereby enhancing the diversity and information content of the training data. Several methods with different generation strategies were designed to address the single-source domain adaptation problem in computer vision tasks. For example, RandConv [26] employed random convolutions for data augmentation. MixStyle integrated style information from instances of randomly selected different domains.
However, previous methods assumed that the target domain shares a similar distribution with the source domain. However, these methods have limitations when applied to scenarios with large distribution gaps, such as medical images and remote sensing images. In contrast with the above data augmentation methods, we use the foundation model to ensure rich prior knowledge instead of generating many images in the source domain, saving computing resources and increasing computation efficiency.
@@ II-B Few Shot Segmentation
Few-shot segmentation (FSS) tasks aim to segment new semantic objects through a limited number of available labelled or unlabeled images that are semantically distinct. Current methods primarily focused on improvements during the meta-learning phase. Prototype-based methods [27, 28] utilized a technique in which representative foreground or background prototypes were extracted from the support data and various strategies were employed to facilitate interactions either between different prototypes or between prototypes and query features. Relation-based methods [29, 30, 31] also succeeded in few-shot segmentation. HSNet [32] built a high correlation using multi-scale dense matching and captures contextual information using 4D convolution. RePRI [33] introduced transductive inference of base class feature extraction that did not require meta-learning. Besides, to apply the generalization to new classes, Lu et al. [4] proposed a Prediction Calibration Network (PCN) for Generalized Few-shot Semantic Segmentation (GFSS), which used a Transformer-based calibration module and cross-attention to reduce class bias and improve segmentation. Chen et al. Moreover, [34] proposed a dual-branch learning method for few-shot semantic segmentation, addressing intra-class and inter-class challenges by enhancing feature representations and generalizability to novel classes. However, these methods primarily focused on segmenting new categories from the same domain with computationally intensive similarity calculation. Due to the significant differences in cross-domain distributions, they failed to be extended to unseen domains.
@@ II-C Cross-domain Few-shot Segmentation
Existing Cross-Domain Few-Shot Learning (CDFSL) methods aim to generalize models to new domains and unseen classes but typically require access to source domain data during pre-training. To reduce the reliance on source domain data, Xu et al. [35] proposed an IM-DCL method for Source-Free Cross-Domain Few-Shot Learning (SF-CDFSL), addressing limited labeled target samples and domain disparities through transductive learning and contrastive learning. However, when dealing with more fine-grained tasks such as segmentation, traditional methods often underperform. In this context, Cross-Domain Few-Shot Segmentation (CD-FSS) has emerged as a specialized area, addressing these challenges with benchmarks and novel strategies.
@@
There are four benchmarks [6] available for CD-FSS standard evaluation. For the ChestX dataset, the image format has been changed from RGB to gray, with a large gap from the original domain. The other two datasets have more edge information requiring high-quality semantic segmentation. RD [36] introduced a novel domain enhancement strategy leveraging a memory mechanism. This approach involved continuously storing domain-style information from the source domain during the training phase. Subsequently, during testing, this stored source information was utilized to enhance the segmentation performance. During testing, source domain information stored in memory was loaded for the target domain feature enhancement. RD [36] offered a direct approach to reduce domain differences and was validated on typical partitioned datasets. For semantic segmentation tasks in autonomous driving applications, PixDA [37] introduced an innovative pixel-by-pixel domain adversarial loss based on three key criteria: (i) aligning the source and target domains for each pixel, (ii) preventing negative transfer on correctly represented pixels, and (iii) regularizing the training of infrequent classes to mitigate overfitting. CDTF [38] achieved cross-domain few-shot segmentation by aligning support and query prototypes. This alignment was realized using an uncertainty-aware contrastive loss and supplemented with a supervised cross-entropy loss and an unsupervised boundary loss as regularization terms. CDTF [38] enabled the model to generalize from the base model to the target domain without requiring additional labels. CD-FSS [39] presented a cross-domain few-shot segmentation framework that leveraged learning from natural domains to assist in rare-disease skin lesion segmentation. This approach was particularly valuable when dealing with limited data for common diseases in the target domain. PBAL [40] came up with a prototype learning and learning technology, which introduced prototype learning and prototype self-training to achieve optimal inter-domain vision and adaptation. However, these methods require a large amount of data for training to achieve a robust model.
@@
Putting aside the previous simple prototype methods, we combine high-level prototype representation with a foundation model, SAM. Besides, we propose a dual prototype matching method for the foreground and background, ensuring fine-grained feature representation.
II-D SAM based Methods
Existing SAM-based segmentation methods incorporate various strategies, including prompt optimization, memory bank feature matching, and adapter modules, to enhance model performance and generalization across domains. Decoupled SAM (DeSAM) is proposed to address the domain shift issue in medical image segmentation [41]. It introduces a prompt-relevant IoU module (PRIM) and a prompt-decoupled mask module (PDMM) to reduce performance degradation caused by poor prompts, achieving enhanced cross-domain robustness on prostate and abdominal datasets. A source domain prior-assisted module is proposed to enhance the generalization of SAM-based medical image segmentation across domains [42]. By utilizing a memory bank to store source domain features, the model matches target domain features with these priors to adapt and improve segmentation accuracy. The CDSG-SAM pipeline is proposed to improve cross-domain few-shot brain tumor segmentation, integrating SAM with a Cross-domain Self-attention (CDS) Adapter and a Self-Generating (SG) Prompt module [43].
III Methology
In this section, we introduce our proposed framework TAVP. First, we describe the problem definition. Then, we present an overview of the proposed approach, followed by a detailed explanation of each technical component of the method.